OpenStudy (calculusxy):

MEDAL! Find the next three terms in each sequence. 1, -3, 9, -27, 81...

OpenStudy (calculusxy):

Can someone help me derive the recursive formula for this sequence as well to find the next three terms?

OpenStudy (xguardians):

What is the pattern that you see?

OpenStudy (calculusxy):

I know that to get to the next term you need to multiply by -3.

OpenStudy (xguardians):

Alright, well now you know how to find the next 3 terms.

OpenStudy (calculusxy):

But I want to find the recursive formula for this problem to help me easily get to the solution.

OpenStudy (xguardians):

As for the recursive formula, it looks something like this: An = -3^n-1

OpenStudy (calculusxy):

How did you get that?

OpenStudy (xguardians):

-3 is what each value changes by, so when you plug the first term, you get. An = -3^(1-1) An = -3^0 = 1 For the second An = -3^(2-1) An -3^1 = -3 etc. etc. It's a geometric sequence.

OpenStudy (xguardians):

An = n^(n-1), where n is the amount it changes each time.

OpenStudy (xguardians):

*the first N.

OpenStudy (calculusxy):

So how did you know that you needed to make (n - 1) as an exponent to the nth term?

OpenStudy (xguardians):

That's the equation for a geometric recursive formula?

OpenStudy (xguardians):

They will always have (n-1)

OpenStudy (calculusxy):

Sorry I didn't learn about geometric recursive formula yet. But in those formulas, will (n-1) always be raised as a power?

OpenStudy (xguardians):

Yes, for geometric ones it will always be that way.

OpenStudy (calculusxy):

Thank you!